above the en the liquid eliquid water held at a temperature of 1.4°C was measured to be temperature was raised to be 55.1°C the recorded total pressure of The pressure of the gas above liquid water gas above the liq of kPa at 55.1°C? (8 points) sured to be 133.6 kPa. What was the vapor pressure of water in units
Answer #1

Assuming the gas is ideal and the volume of the container doesn't change then we can write \frac {P_{1}}{T_{1}}=\frac {P_{2}}{T_{2}}

P1 and P2 are initial and final temperature respectively and T1 and T2 are initial and final temperatures

then We know P1=100.2 and P2=x, T1=1.4+273=274.4K and T2=273+55.1=328.1K

Hence P2=(T2*P1)/T1=(328.1*100.2)/274.4=119.8 kPa

hence The remaining pressure out of 133.6 kPa is contributed by the vapor pressure of the water.

Hence VP of water at 55.1 C = 133.6-119.8=13.79 kPa

Please do give this answer a thumbs up as it matters a lot and leave a comment in case there is any doubt.